# Connection between galaxy morphology and dark-matter halo structure II: predicting disk structure from dark-matter halo properties

### v1

@@ 5 Conclusion
Across all targets, RF and SR consistently identify concentration, spin, mass, and assembly history as the most predictive features. Despite lower accuracy, DMO-based predictions broadly agree with hydro-based trends, suggesting that many disk-halo relations are robust to baryonic differences. The SR relations presented here outperform previous analytic prescriptions (Kravtsov13; Mo98; Jiang19).
@@ A.1 Random Forest
We use the RF regressor from scikit-learn (Pedregosa11). The hyperparameters tuned in the module are listed below while other hyperparameters remain default settings:
@@ A.2 Shapley Additive Explanations Values
To interpret regression results from the complex tree structures in RF, we use SHAP (Lundberg17), a game-theoretic method. SHAP evaluates how much each feature contributes to a prediction by comparing the model’s output with and without that feature across many possible feature combinations. A SHAP value is the weighted average of these marginal contributions, capturing both the direction and strength of a feature’s effect.
@@ A.3 Symbolic Regression
For finding empirical and interpretable equations from the most important features identified by SHAP, we apply the SR (e.g. Koza92) equation search algorithm PySR (Cranmer23).
@@
The loss function $\mathcal{L}$ used in this work combines a normalized mean squared error term $\mathcal{L}_{\rm NMSE}$ with a normalized quantile loss term $\mathcal{L}_{\rm NQL}$. The normalized mean squared error term takes form
$\mathcal{L}_{\rm NMSE}=(y-\hat{y})^{2}/y^{2}$ (A1)
where $y$ is the real values and $\hat{y}$ is the predicted values. The normalized quantile loss captures the asymmetric penalties for underestimation and overestimation across multiple quantiles $\tau_{i}$ which taken from [0.1,0.25,0.5,0.75,0.9]. The quantile loss is computed as
$\quad+(1-\tau_{i})\max\Big(-\frac{y-\hat{y}}{y},0\Big)\Bigg]$ (A2)
The final loss function is taken from $\mathcal{L}=\mathcal{L}_{\rm NMSE}+\mathcal{L}_{\rm NQL}$.
Besides the equations, PySR also returns the loss $\mathcal{L}$ and complexity $\mathcal{C}$ for each equation. More complex equations typically achieve lower loss but are harder to interpret and implement into semi-analytical models. To balance complexity and loss, we designed a score to automatically select equations from all possible solutions. We first construct normalized loss $\tilde{\mathcal{L}}$ and normalized complexity $\tilde{\mathcal{C}}$ by
$\tilde{\mathcal{L}}=\frac{\mathcal{L}-\min(\mathcal{L})}{\max(\mathcal{L})-\min(\mathcal{L})},\ \tilde{\mathcal{C}}=\frac{\mathcal{C}-\min(\mathcal{C})}{\max(\mathcal{C})-\min(\mathcal{C})}$ (A3)
Then the score is defined as
$\text{Score}=\Bigg[w\log(1+\tilde{\mathcal{L}})+(1-w)\log(1+\tilde{\mathcal{C}})\Bigg]^{-1}$ (A4)
where $w$ is the weight. This formula first increases with complexity, reaching a maximum, and then decreases with complexity. We choose the equation at the maximum point. A higher $w$ results in greater complexity but lower loss at the maximum point. We set the $w$ to 0.7 which leads a equation with complexity $\sim 50$.
@@ C.3 Mass fraction
Appendix D Bias in sample selection
Following Paper I, we require a minimum of 1000 stellar and 1000 DM particles. The stellar threshold ensures robust morphological decomposition but is the more restrictive condition. Since galaxies with 1,000 stellar particles typically inhabit halos well above the mass equivalent of 1000 DM particles, our sample and the measurements for halo quantities becomes incomplete at the low-mass end. From a physical perspective, we ignore galaxies that are strongly DM-dominated, which may represent a substantial fraction of the population.
Such incompleteness biases the comparisons between predictions and data. Note that because predictions are calculated from the SR relation using halo measurements, they effectively act as halo properties. This bias is evident in three perspectives. First, while binning by predictions yields medians close to the data, the reverse (binning by data) exposes a bias at the low-mass end. Second, the visual agreement and $R^{2}$ values improve when plotting physical quantities (e.g., $M_{\star}$) rather than normalized ones (e.g., $M_{\star}/M_{\rm vir}$). Third, the maximum halo mass for the galaxies with 1000 stellar particles is $\sim 10^{10.8}M_{\odot}$. If one test the RF and SR using samples with halo mass above this threshold, the performance will be better. We therefore caution the use of SR relations for samples below this threshold.

### v2

@@ V Conclusion
Across all targets, RF and SR consistently identify concentration, spin, mass, and assembly history as the most predictive features. Despite lower accuracy, DMO-based predictions broadly agree with hydro-based trends, suggesting that disk–halo relations derived from DMO runs still possess decent predictive power. The SR relations presented here outperform previous analytic prescriptions (Kravtsov, 2013; Mo et al., 1998; Jiang et al., 2019).
@@ A.1 Random Forest
We use the RF regressor from scikit-learn (Pedregosa et al., 2011). The hyperparameters tuned in the module are listed below while other hyperparameters remain default settings:
@@ A.2 Shapley Additive Explanations Values
To interpret regression results from the complex tree structures in RF, we use SHAP (Lundberg and Lee, 2017), a game-theoretic method. SHAP evaluates how much each feature contributes to a prediction by comparing the model’s output with and without that feature across many possible feature combinations. A SHAP value is the weighted average of these marginal contributions, capturing both the direction and strength of a feature’s effect.
@@ A.3 Symbolic Regression
For finding empirical and interpretable equations from the most important features identified by SHAP, we apply the SR (e.g. Koza, 1992) equation search algorithm PySR (Cranmer, 2023).